# Scale Where It Matters: Training-Free Localized Scaling for Diffusion Models

### v2

@@ 1 Introduction
Diffusion models have become the de-facto standard for high-quality image generation, owing to their strong scalability with data, model size, and compute. This scalability has driven remarkable advances in text-to-image synthesis [ho2020denoising, saharia2022photorealistic, rombach2022high, ruiz2023dreambooth, sun2024coma, sun2025ouroboros], establishing scaling laws as a guiding principle for building more capable models. While most prior work has focused on scaling at training time, recent studies show that test-time scaling (TTS), allocating additional compute during inference, can also significantly improve sample quality and overall performance [nichol2021glide, esser2024scaling, peebles2023scalable, ma2025scaling, liu2024correcting]. Despite its promise, existing TTS research remains limited in scope, leaving open questions about how to use inference-time compute more effectively.
@@
Existing TTS methods can be broadly grouped into three categories. The first is denoising step scaling, which improves quality by increasing the number of sampling steps [song2020score, lu2022dpm]. However, these gains saturate quickly and plateau around 50 steps, with further increases offering negligible benefit. The second is Best-of-$N$ search, where generates $N$ samples and selects the best one by verifier [liu2024correcting, wang2023exploring]. While simple, this brute-force approach treats each candidate as an independent sample from scratch, overlooking the fact that even imperfect images may be substantially improved through local corrections. As a result, potentially promising samples are discarded, and computation is wasted on redundant global search. The third is trajectory/noise search, which perturbs the initial noise or explores alternative sampling paths [xu2023restart, ramesh2025test]. Although more fine-grained, it still operates at the full-image level. Searching over all regions can inadvertently disturb areas that are already of high quality, thereby leading to inefficiency and instability. Despite their differences, all three categories share a fundamental limitation: they operate at the full-image level, as shown in Fig. 1. Consequently, they overlook the inherent spatial heterogeneity of image quality and fail to exploit the potential of localized refinement.
@@
This observation naturally motivates an orthogonal direction: localized TTS, where only defective regions are resampled while preserving high-quality content [cao2025temporal], thereby substantially reducing the search space. Yet turning this idea into practice introduces two key challenges. The first is accurate defect localization: since the distribution of artifacts is complex and prompt-dependent, reliably identifying regions that truly require correction is non-trivial [liu2024improved, zhang2023perceptual]. The second is maintaining consistency in local resampling: as the sampling trajectory is globally defined, locally modifying only a subset of regions may introduce incompatibility, leading to semantic drift, stylistic inconsistency, or boundary artifacts that degrade perceptual quality [song2020score, xu2023restart].
In this paper, we propose LoTTS (Prompt-Guided Localized Test-Time Scaling), the first fully training-free localized TTS framework. Specifically, LoTTS consists of two key components: defect localization and consistency maintenance. For defect localization, LoTTS exploits attention signals from diffusion models to construct a quality-aware mask. We contrast semantically equivalent but quality-differentiated prompts (e.g., “a high-quality image of {p}” vs. “a low-quality image of {p}”) and compute the difference between their cross-attention maps, which highlights candidate defective regions [hertz2022prompt, chefer2023attend, chung2024style]. To suppress spurious activations, these difference maps are reweighted with the original prompt attention, filtering out irrelevant background responses. The resulting signals are then propagated via self-attention to enforce spatial coherence, yielding an automated defect mask without reliance on external predictors or manual annotations. For consistency maintenance, LoTTS introduces a localized defect-aware resampling mechanism. After initial sampling, we inject controlled noise into the defective regions of the generated image at an intermediate timestep, followed by localized denoising [meng2021sdedit]. To reconcile local refinements with preserved content, we perform a few global denoising steps, harmonizing style and structure across the image. This ensures that defects are corrected without sacrificing global fidelity. Furthermore, LoTTS is fully plug-and-play and can be seamlessly applied to both diffusion- and flow-based generative models [song2020score, lu2022dpm]. Our main contributions are summarized as follows:
  • We propose LoTTS, the first training-free localized Test-Time Scaling framework, which shifts the focus of inference-time compute from the entire image to defect regions, addressing the inefficiency of global search.
  • We introduce an efficient localized noise search mechanism, comprising prompt-driven quality map generation and localized defect-aware resampling.
@@
Figure 2: Overview of LoTTS. LoTTS begins by generating multiple candidate images from different noise seeds, identifies defective regions via high-/low-quality prompt contrast, constructs a quality-aware defect mask, injects noise selectively within the masked areas, performs localized denoising with spatial and temporal consistency, and finally selects the best refined sample with a verifier. This pipeline selectively improves low-quality regions while preserving global fidelity.
@@ 2 Related work
Image Generation and Evaluation.   Image generation has advanced from GANs [goodfellow2020generative, karras2019style] and autoregressive models [van2016pixel, van2016conditional] to flow-based and diffusion models. GANs achieve high fidelity but are unstable, while autoregressive models capture dependencies yet remain slow. Early flows [kingma2018glow] provide exact likelihoods but scale poorly, whereas recent flow matching and rectified flow methods [lipman2022flow, albergo2023stochastic, kim2025inference] connect closely to diffusion and offer efficient alternatives. Diffusion models [ho2020denoising, rombach2022high] now dominate, delivering state-of-the-art text-to-image synthesis [rombach2022high, podell2023sdxl]. In parallel, evaluation methods have evolved: earlier work often collapsed quality into a single global score [wang2004image, salimans2016improved, heusel2017gans], while spatial indicators relied on texture- or frequency-based heuristics [yu2019attributing, durall2020watch] or classifier- and VLM-based predictors [liu2021aligning, zhang2023perceptual, liu2024improved], typically requiring supervision and external datasets. In contrast, LoTTS leverages the diffusion model’s inherent attention to localize defects automatically, enabling training-free, localized refinement.
@@
Defect Localization.   A substantial body of work aims to localize defects in generated images. Early methods analyze GAN or super-resolution artifacts through spatial or frequency cues [yu2019attributing, durall2020watch, dzanic2020fourier, you2025uncovering]. More recent approaches train supervised models to predict perceptual-error or artifact heatmaps [zhang2023perceptual, li2024g, ren2025otsurv], or pursue defect detection via uncertainty estimation [kou2023bayesdiff] and diffusion-trajectory analysis [cao2025temporal]. In contrast to methods that rely on external predictors, supervised training, or heavy post-hoc analysis, LoTTS derives defect localization directly from the diffusion model itself. Our masks are computed from attention responses to positive and negative prompts, leveraging the model’s inherent cross-modal alignment. This makes LoTTS fully training-free and inherently consistent with the generative semantics of the T2I model.
@@
Generation Quality Enhancement.   Image Editing in modern T2I systems is typically implemented as a post-processing stage. The cascade paradigm, consisting of a generator followed by a refiner, has become a standard design, as in IF [genIF, you2022megan], SDXL [podell2023sdxl] and SD Cascade [pernias2023wurstchen], where an additional pass refines the entire image to enhance global fidelity. However, such refiners usually require separate training and introduce non-negligible computational overhead. In contrast, SDEdit [meng2021sdedit] shows that diffusion models can perform localized edits via noise–denoise updates within user-specified masks, suggesting the potential of training-free localized refinement. This localized resampling mechanism inspires our approach, which extends SDEdit’s manual editing to automated quality-aware refinement at test time.
@@
Test-Time Scaling in Vision.   Some test-time scaling methods have been proposed to enhance diffusion model generation by allocating more computation at inference. Early work simply increased denoising steps, but improvements saturate quickly beyond a certain number of function evaluations (NFE) [karras2022elucidating, song2020denoising, song2020score]. Recent studies therefore explore alternative directions, such as Best-of-$N$ search, where multiple candidates are generated from different noise seeds and a verifier selects the best one [liu2024correcting, wang2023exploring], searching over noise initializations [song2020score, xu2023restart], optimizing sampling trajectories [ramesh2025test] with verifier feedback [song2020denoising, karras2022elucidating, liu2022pseudo, lu2022dpm, salimans2022progressive], or adopting evolutionary [he2025scaling] and tree-search methods [yoon2025fast]. Unlike these approaches, which all operate on the entire image and require full regeneration, our LoTTS performs localized TTS by concentrating on low-quality regions for greater efficiency.
3 Preliminary
@@
Diffusion Models.   Diffusion models transform a simple source distribution, e.g.a standard Gaussian, into a target data distribution $p_{0}$. In diffusion models [sohl2015deep, ho2020denoising], the forward process gradually corrupts clean data with noise, as
@@
where $\alpha(t)$ and $\sigma(t)$ denote the noise schedule, and $t\in[0,T]$. To recover data from its diffused representation, diffusion models generally rely on Stochastic Differential Equation(SDE)-based sampling during inference [song2020score, song2020denoising], which introduces stochasticity at every denoising step:
$\bm{x}_{t-1}$ $=\sqrt{\alpha(t-1)}\left(\frac{\bm{x}_{t}-\sqrt{1-\alpha(t)}\,\epsilon_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha(t)}}\right)$
$\quad+\sqrt{1-\alpha(t-1)-\sigma(t)^{2}}\,\epsilon_{\theta}(\bm{x}_{t},t)$
$\quad+\sigma(t)\,\epsilon(t).$ (2)
Here, $\epsilon_{\theta}(\bm{x}_{t},t)$ denotes the predicted noise at step $t$, $\alpha(t)$ and $\sigma(t)$ are the noise schedule parameters, and $\epsilon_{t}\sim\mathcal{N}(0,I)$ is a standard Gaussian noise.
Flow Models.   Flow models [lipman2022flow, albergo2023stochastic] parameterize the velocity field $u_{t}\in\mathbb{R}^{d}$ and generate samples by solving the Flow Ordinary Differential Equation (ODE) [song2020score] backward from $t=T$ to $t=0$:
@@
where $u_{t}(\bm{x}_{t})$ denotes the velocity field learned by the flow model, and $dt$ is an infinitesimal step along the reverse-time ODE. This deterministic dynamics evolves $\bm{x}_{t}$ continuously in time, producing identical outcomes for the same input and limiting the applicability of test-time scaling methods [kim2025inference], which require stochasticity to explore diverse trajectories. To address this limitation, recent studies propose that the deterministic Flow-ODE could be reformulated into an equivalent SDE [albergo2023stochastic, ma2024sit, patel2024exploring, kim2025inference, singh2024stochastic]. The resulting stochastic process can be written as:
@@
where the score function $\nabla\log p_{t}(\bm{x}_{t})$ can be estimated from the velocity field $u_{t}$ [singh2024stochastic]), and the Brownian motion term $d\bm{w}$ introduces stochasticity at each sampling step. This enables LoTTS to naturally apply to Flow-Models.
@@ 4 Method
4.1 Preliminaries
Motivation and Challenges.   Existing test-time scaling methods, such as Best-of-$N$ or trajectory search, apply uniform resampling to the entire image, overlooking spatial heterogeneity in quality and thereby wasting computation or even degrading well-formed regions. This observation motivates the consideration of a more targeted form of TTS, one that selectively resamples only defective regions while preserving high-quality areas. By reducing the effective search space and avoiding unnecessary perturbation to clean regions, such a localized TTS paradigm could potentially improve both efficiency and stability compared with full-image resampling. However, realizing this paradigm raises two key challenges: (1) Defect Localization: how to generate reliable resample masks that identify low-quality regions, and (2) Consistency Maintenance: how to perform localized resampling within these regions without disrupting the rest of the image.
@@ 4.1 Preliminaries
For completeness, pseudocode is given in Appendix B, and Appendix G provides a theoretical analysis showing that LoTTS can be proven to achieve a higher expected quality gain under stated conditions.
@@ 4.2 Defect Localization
A key challenge in localized test-time scaling is to identify which regions of a generated image truly require refinement. Since image degradations are usually local and diverse, a reliable defect localization mechanism is indispensable. Existing verifiers typically provide only global quality scores without spatial resolution, making them unsuitable for localized resampling. We leverage the intrinsic attention signals of diffusion models to automatically infer defect regions. The core intuition is that cross-attention with a “low-quality“ prompt points to artifact-prone regions, and comparing it with a “high-quality“ prompt makes those local defects stand out. To obtain stable and semantically meaningful masks, we further propagate and reconcile these information flows.
As illustrated in Fig. 3, the process has four stages: (1) Prompt-driven Discrimination: compute cross-attention maps under high-/low-quality prompts and take their difference to highlight candidate defect regions. (2) Context-aware Propagation: refine these signals by propagating across spatially similar positions, mitigating noise, and enforcing local coherence. (3) Semantic-guided Reweighting: combine with original prompt attention to suppress irrelevant background. (4) Quality-aware Mask Generation: binarize by percentile to control the resampling ratio.
Figure 3: LoTTS Framework for Defect Localization. The pipeline consists of four stages: prompt-driven discrimination, context-aware propagation, semantic-guided reweighting, and quality-aware mask generation.
Prompt-driven Discrimination.   The intuition is that cross-attention maps reflect how different prompts attend to image regions: a “low-quality” prompt tends to focus on artifact-prone areas, while a “high-quality” prompt attends to cleaner regions. By contrasting the two, we can highlight potential defects.
Following the common setting [liu2024towards, helbling2025conceptattention], we obtain a spatial cross-attention vector $\mathrm{CA}\in\mathbb{R}^{S}$ by averaging over tokens, heads, and selected layers, where $S=H_{s}\times W_{s}$ denotes the number of spatial positions. For a given prompt $p$, we construct three variants: a positive prompt (“A high-quality image of {pp}”), a negative prompt (“A low-quality image of {pp}”), and the origin prompt (the original $p$). This yields three attention vectors $\mathrm{CA}_{\mathrm{pos}},\mathrm{CA}_{\mathrm{neg}},\mathrm{CA}_{\mathrm{orig}}\in\mathbb{R}^{S}$. We then define a contrastive cross-attention map as:
@@
which emphasizes spatial locations where the “bad” prompt receives higher attention than the “good” prompt. The origin vector $\mathrm{CA}_{\mathrm{orig}}$ will later serve as a foreground prior in our aggregation step.
@@
$\mathrm{CA}^{\prime}=\mathrm{SA}\times\mathrm{CA},\qquad\mathrm{SA}=\mathrm{Softmax}{(\frac{QQ^{\top}}{\sqrt{d}})}.$ (6)
@@
Semantic-guided Reweighting.   Although the contrastive map $\mathrm{CA}^{\prime}_{\mathrm{diff}}$ can reveal defect-prone areas, it often assigns high scores to background regions with little semantic content (e.g., large sky areas). The intuition is that truly meaningful defects should also lie within the semantic foreground defined by the original prompt. To encode this prior, we combine the contrastive map with the original attention map $\mathrm{CA}^{\prime}_{\mathrm{orig}}$, which serves as a soft foreground mask:
@@
Quality-Aware Mask Generation.   Finally, we need to convert the aggregated quality map into a binary mask that specifies which regions should be resampled. The key idea is to only refine the most degraded areas rather than the entire image. To achieve this, we keep the top $r$ proportion of spatial positions with the highest defect scores, ensuring that resampling is both targeted and controllable. Formally, the mask $M\in\{0,1\}^{S}$ is given by:
@@
where $\mathrm{Perc}(P,\,1-r)$ is the (1r)(1-r)-quantile of the quality map $P$, and $r\in(0,1)$ controls the fraction of resampled regions. The mask is then reshaped into the spatial grid $\mathbf{M}\in\{0,1\}^{H_{s}\times W_{s}}$ to guide localized refinement. We compute the attention mask at $t=0$, matching the final image.
Table 1: Quantitative results on three benchmarks (Pick-a-Pic, DrawBench, COCO2014). Across human-aligned metrics (HPS, AES, Pick, IR) and automated metrics (FID, CLIP), LoTTS consistently outperforms Resampling, Best-of-$N$, and Particle Sampling baselines.

### v3

@@ 1 Introduction
Diffusion models have become the de-facto standard for high-quality image generation, owing to their strong scalability with data, model size, and compute. This scalability has driven remarkable advances in text-to-image synthesis (ho2020denoising; saharia2022photorealistic; rombach2022high; ruiz2023dreambooth; sun2024coma; sun2025ouroboros), establishing scaling laws as a guiding principle for building more capable models. While most prior work has focused on scaling at training time, recent studies show that test-time scaling (TTS), allocating additional compute during inference, can also significantly improve sample quality and overall performance (nichol2021glide; esser2024scaling; peebles2023scalable; ma2025scaling; liu2024correcting). Despite its promise, existing TTS research remains limited in scope, leaving open questions about how to use inference-time compute more effectively.
@@
Existing TTS methods can be broadly grouped into three categories. The first is denoising step scaling, which improves quality by increasing the number of sampling steps (song2020score; lu2022dpm). However, these gains saturate quickly and plateau around 50 steps, with further increases offering negligible benefit. The second is Best-of-$N$ search, which generates $N$ samples and selects the best one via a verifier (liu2024correcting; wang2023exploring). While simple, this brute-force approach treats each candidate as an independent sample from scratch, overlooking the fact that even imperfect images may be substantially improved through local corrections. As a result, potentially promising samples are discarded, and computation is wasted on redundant global search. The third is trajectory/noise search, which perturbs the initial noise or explores alternative sampling paths (xu2023restart; ramesh2025test). Although more fine-grained, searching over all regions can inadvertently disturb areas that are already of high quality, leading to inefficiency and instability. Despite their differences, all three categories share a fundamental limitation: they operate at the full-image level, as shown in Fig. 1. Consequently, they overlook the inherent spatial heterogeneity of image quality and fail to exploit the potential of localized refinement.
@@
This observation naturally motivates an orthogonal direction: localized TTS, where only defective regions are resampled while preserving high-quality content (cao2025temporal), thereby substantially reducing the search space. Yet turning this idea into practice introduces two key challenges. The first is accurate defect localization: since the distribution of artifacts is complex and prompt-dependent, reliably identifying regions that truly require correction is non-trivial (liu2024improved; zhang2023perceptual). The second is maintaining consistency in local resampling: as the sampling trajectory is globally defined, locally modifying only a subset of regions may introduce incompatibility, leading to semantic drift, stylistic inconsistency, or boundary artifacts that degrade perceptual quality (song2020score; xu2023restart).
Our key observation is that diffusion models already encode spatially resolved quality information in their cross-attention maps: by contrasting attention under quality-differentiated prompts (e.g. “high-quality” vs. “low-quality”), defective regions naturally stand out – enabling zero-shot defect localization without any external supervision. Building on this insight, we propose LoTTS (Prompt-Guided Localized Test-Time Scaling), the first fully training-free localized TTS framework that addresses both challenges in a unified pipeline. For defect localization, LoTTS contrasts cross-attention maps under quality-differentiated prompts to produce a quality-aware mask that highlights artifact-prone regions, without relying on any external predictors or annotations. For consistency maintenance, LoTTS injects noise only into the masked regions at an intermediate timestep and performs localized denoising (meng2021sdedit), followed by a global harmonization pass to ensure seamless integration. The framework is fully plug-and-play and applies to both diffusion- and flow-based models (song2020score; lu2022dpm). Our main contributions are summarized as follows:
  • We propose LoTTS, the first training-free localized Test-Time Scaling framework. Its core is a prompt-driven defect localization mechanism that contrasts cross-attention signals to produce quality-aware masks without any external supervision, paired with a consistency-preserving resampling strategy that refines only the masked regions while maintaining coherence with the rest of the image.
  • We provide theoretical analysis showing that localized resampling achieves a higher expected quality gain than global resampling under stated conditions, formally justifying the localized TTS paradigm.
@@
Figure 2: Overview of LoTTS. Given a text prompt, LoTTS first generates candidate images from different noise seeds. It then localizes defective regions using high-/low-quality prompt contrast and constructs a quality-aware mask. Noise is injected only inside the masked regions, followed by localized denoising with spatial and temporal consistency. A verifier finally selects the best refined sample.
@@ 2 Related Work
Image Generation and Evaluation. Image generation has advanced from GANs (goodfellow2020generative; karras2019style) and autoregressive models (van2016pixel; van2016conditional) to flow-based and diffusion models. GANs achieve high fidelity but are unstable, while autoregressive models capture dependencies yet remain slow. Early flows (kingma2018glow) provide exact likelihoods but scale poorly, whereas recent flow matching and rectified flow methods (lipman2022flow; albergo2023stochastic; kim2025inference) connect closely to diffusion and offer efficient alternatives. Diffusion models (ho2020denoising; rombach2022high) now dominate, delivering state-of-the-art text-to-image synthesis (rombach2022high; podell2023sdxl). In parallel, evaluation methods have evolved: earlier work often collapsed quality into a single global score (wang2004image; salimans2016improved; heusel2017gans), while spatial indicators relied on texture- or frequency-based heuristics (yu2019attributing; durall2020watch) or classifier- and VLM-based predictors (liu2021aligning; zhang2023perceptual; liu2024improved), typically requiring supervision and external datasets. In contrast, LoTTS leverages the diffusion model’s inherent attention to localize defects automatically, enabling training-free, localized refinement.
@@
Defect Localization. A substantial body of work aims to localize defects in generated images. Early methods analyze GAN or super-resolution artifacts through spatial or frequency cues (yu2019attributing; durall2020watch; dzanic2020fourier; you2025uncovering). More recent approaches train supervised models to predict perceptual-error or artifact heatmaps (zhang2023perceptual; li2024g; ren2025otsurv), or pursue defect detection via uncertainty estimation (kou2023bayesdiff; liu2025together; zhang2025supervise) and diffusion-trajectory analysis (cao2025temporal). In contrast to methods that rely on external predictors, supervised training, or heavy post-hoc analysis, LoTTS derives defect localization directly from the diffusion model itself. Our masks are computed from attention responses to positive and negative prompts, leveraging the model’s inherent cross-modal alignment. This makes LoTTS fully training-free and inherently consistent with the generative semantics of the T2I model.
@@
Generation Quality Enhancement. Image editing in modern T2I systems is typically implemented as a post-processing stage. The cascade paradigm, consisting of a generator followed by a refiner, has become a standard design, as in IF (genIF; you2022megan), SDXL (podell2023sdxl) and SD Cascade (pernias2023wurstchen), where an additional pass refines the entire image to enhance global fidelity. However, such refiners usually require separate training and introduce non-negligible computational overhead. In contrast, SDEdit (meng2021sdedit) shows that diffusion models can perform localized edits via noise–denoise updates within user-specified masks, suggesting the potential of training-free localized refinement. This localized resampling mechanism inspires our approach, which extends SDEdit’s manual editing to automated quality-aware refinement at test time.
@@
Test-Time Scaling in Vision. Some test-time scaling methods have been proposed to enhance diffusion model generation by allocating more computation at inference. Early work simply increased denoising steps, but improvements saturate quickly beyond a certain number of function evaluations (NFE) (karras2022elucidating; song2020denoising; song2020score). Recent studies therefore explore alternative directions, such as Best-of-$N$ search, where multiple candidates are generated from different noise seeds and a verifier selects the best one (liu2024correcting; wang2023exploring), searching over noise initializations (song2020score; xu2023restart), optimizing sampling trajectories (ramesh2025test) with verifier feedback (song2020denoising; karras2022elucidating; liu2022pseudo; lu2022dpm; salimans2022progressive; you2024calibrating), or adopting evolutionary (he2025scaling) and tree-search methods (yoon2025fast). Unlike these approaches, which all operate on the entire image and require full regeneration, our LoTTS performs localized TTS by concentrating on low-quality regions for greater efficiency.
3 Preliminaries
@@
Diffusion Models. Diffusion models transform a simple source distribution, e.g. a standard Gaussian, into a target data distribution $p_{0}$. In diffusion models (sohl2015deep; ho2020denoising), the forward process gradually corrupts clean data with noise, as:
@@
where $\alpha(t)$ and $\sigma(t)$ denote the noise schedule, and $t\in[0,T]$. To recover data from its diffused representation, diffusion models generally rely on Stochastic Differential Equation (SDE)-based sampling during inference (song2020score; song2020denoising), which introduces stochasticity at every denoising step:
$\bm{x}_{t-1}=\alpha(t{-}1)\,\hat{\bm{x}}_{0}+\sqrt{\sigma(t{-}1)^{2}-\tilde{\sigma}_{t}^{2}}\,\epsilon_{\theta}(\bm{x}_{t},t)+\tilde{\sigma}_{t}\,\epsilon_{t},$ (2)
where $\hat{\bm{x}}_{0}=(\bm{x}_{t}-\sigma(t)\,\epsilon_{\theta}(\bm{x}_{t},t))/{\alpha(t)}$ is the predicted clean image, $\tilde{\sigma}_{t}$ controls the stochasticity of sampling, and $\epsilon_{t}\sim\mathcal{N}(0,I)$.
Flow Models. Flow models (lipman2022flow; albergo2023stochastic) parameterize the velocity field $u_{t}\in\mathbb{R}^{d}$ and generate samples by solving the Flow Ordinary Differential Equation (ODE) (song2020score) backward from $t=T$ to $t=0$:
@@
where $u_{t}(\bm{x}_{t})$ denotes the velocity field learned by the flow model, and $dt$ is an infinitesimal step along the reverse-time ODE. This deterministic dynamics evolves $\bm{x}_{t}$ continuously in time, producing identical outcomes for the same input and limiting the applicability of test-time scaling methods (kim2025inference), which require stochasticity to explore diverse trajectories. To address this limitation, recent studies propose that the deterministic Flow-ODE could be reformulated into an equivalent SDE (albergo2023stochastic; ma2024sit; patel2024exploring; kim2025inference; singh2024stochastic). The resulting stochastic process can be written as:
@@
where the score function $\nabla\log p_{t}(\bm{x}_{t})$ can be estimated from the velocity field $u_{t}$ (singh2024stochastic), and the Brownian motion term $d\bm{w}$ introduces stochasticity at each sampling step. This enables LoTTS to naturally apply to flow models.
@@ 4 Method
4.1 Overview
Motivation and Challenges. Existing test-time scaling methods, such as Best-of-$N$ or trajectory search, apply uniform resampling to the entire image, overlooking spatial heterogeneity in quality and thereby wasting computation or even degrading well-formed regions. This motivates a more targeted form of TTS that selectively resamples only defective regions, reducing the search space and improving both efficiency and stability. However, realizing this paradigm raises two key challenges: (1) Defect Localization: how to generate reliable resample masks that identify low-quality regions, and (2) Consistency Maintenance: how to perform localized resampling within these regions without disrupting the rest of the image.
@@ 4.1 Overview
Pseudocode is given in Appendix B, and Appendix G proves that localized resampling achieves a higher expected quality gain than global resampling under mild conditions.
@@ 4.2 Defect Localization
We leverage the intrinsic attention signals of diffusion models to localize defects: cross-attention with a “low-quality” prompt highlights artifact-prone regions, and contrasting it with a “high-quality” prompt makes those defects stand out. To obtain stable and semantically meaningful masks, we further propagate and reconcile these signals.
As illustrated in Fig. 3, starting from extracted cross-attention maps, the process consists of four stages: (1) Prompt-driven Discrimination: take the difference between high-/low-quality cross-attention maps to highlight candidate defect regions. (2) Context-aware Propagation: refine these signals by propagating across spatially similar positions, mitigating noise, and enforcing local coherence. (3) Semantic-guided Reweighting: combine with original prompt attention to suppress irrelevant background. (4) Quality-aware Mask Generation: binarize by percentile to control the resampling ratio.
Figure 3: Defect Localization Overview. LoTTS localizes defects from cross-attention maps through four stages: prompt-driven discrimination, context-aware propagation, semantic-guided reweighting, and quality-aware mask generation.
Prompt-driven Discrimination. Cross-attention maps reflect how different prompts attend to image regions: a “low-quality” prompt tends to focus on artifact-prone areas, while a “high-quality” prompt attends to cleaner regions. By contrasting the two, we can highlight potential defects.
Following the common setting (liu2024towards; helbling2025conceptattention), we obtain a spatial cross-attention vector $\mathrm{CA}\in\mathbb{R}^{S}$ by averaging over tokens, heads, and selected layers, where $S=H_{s}\times W_{s}$ denotes the number of spatial positions. For a given prompt $p$, we construct three variants: a positive prompt (“A high-quality image of {pp}”), a negative prompt (“A low-quality image of {pp}”), and the origin prompt (the original $p$). This yields three attention vectors $\mathrm{CA}_{\mathrm{pos}},\mathrm{CA}_{\mathrm{neg}},\mathrm{CA}_{\mathrm{orig}}\in\mathbb{R}^{S}$. We then define a contrastive cross-attention map as:
@@
which emphasizes spatial locations where the negative prompt receives higher attention than the positive prompt. The origin vector $\mathrm{CA}_{\mathrm{orig}}$ will later serve as a foreground prior in our aggregation step.
@@
$\mathrm{CA}^{\prime}=\mathrm{SA}\times\mathrm{CA},\qquad\mathrm{SA}=\mathrm{Softmax}\!\left(\frac{QQ^{\top}}{\sqrt{d}}\right).$ (6)
@@
Semantic-guided Reweighting. Although the contrastive map $\mathrm{CA}^{\prime}_{\mathrm{diff}}$ can reveal defect-prone areas, it often assigns high scores to background regions with little semantic content (e.g. large sky areas). Since meaningful defects should lie within the semantic foreground, we combine the contrastive map with the original attention map $\mathrm{CA}^{\prime}_{\mathrm{orig}}$, which serves as a soft foreground mask:
@@
Quality-Aware Mask Generation. Finally, we convert the aggregated quality map into a binary mask by keeping the top $r$ proportion of spatial positions with the highest defect scores, ensuring that resampling is both targeted and controllable. Formally, the mask $M\in\{0,1\}^{S}$ is given by:
@@
where $\mathrm{Perc}(P,\,1-r)$ is the (1r)(1-r)-quantile of the quality map $P$, and $r\in(0,1)$ controls the fraction of resampled regions. The mask is then reshaped into the spatial grid $\bm{M}\in\{0,1\}^{H_{s}\times W_{s}}$ to guide localized refinement. We compute the attention mask at $t=0$, matching the final image.
Figure 4: Localized Resampling Process. Initial Denoising: standard denoising from $\bm{x}_{T}$ produces $\bm{x}_{0}$ with localized artifacts (IR = 0.54). Localized Resample: LoTTS injects noise within the defect mask $\bm{M}$ at step $t_{0}$, then performs Masked Refinement followed by Global Integration, yielding $\tilde{\bm{x}}_{0}$ with improved quality (IR = 1.03).
4.3 Consistency Maintenance
With reliable defect masks from the previous step, the next challenge is how to resample the identified regions without disrupting the rest of the image. Restricting updates only to masked regions often introduces boundary artifacts or semantic drift. To address these issues, LoTTS maintains spatial, temporal, and holistic consistency during refinement, as illustrated in Fig. 4.
Spatial Consistency. Resampling in diffusion models typically begins by perturbing the latent representation. If perturbation is applied only within the mask, the noise distribution becomes inconsistent with surrounding regions, creating visible seams. We avoid this by injecting comparable noise into both masked and unmasked areas, which balances noise levels and ensures smooth transitions across boundaries. Formally, given the clean latent $\bm{x}_{0}$ and binary mask $\bm{M}$, we initialize the perturbed latent at timestep $t_{0}$ as:
$\bm{x}_{t_{0}}=\alpha(t_{0})\bm{x}_{0}+\sigma(t_{0})((1-\bm{M})\odot\bm{z}_{\mathrm{bg}}+\bm{M}\odot\bm{z}_{\mathrm{mask}}),$ (9)
where $\bm{z}_{\mathrm{bg}}\!\sim\!\mathcal{N}(0,I)$ and $\bm{z}_{\mathrm{mask}}\!\sim\!\mathcal{N}(0,I)$.
Temporal Consistency. A second challenge is preserving the global semantics of the image. Restarting from pure noise discards structure and forces the model to regenerate the entire scene. Instead, following Meng et al. (meng2021sdedit), we resume denoising from an intermediate step $t_{0}$, so that global content is retained while localized corrections remain possible. The masked reverse update is:
$\begin{split}\bm{x}_{t-\Delta t}=&\,(1-\bm{M})\odot\big(\alpha(t{-}\Delta t)\bm{x}_{0}+\sigma(t{-}\Delta t)\bm{z}\big)\\ &+\bm{M}\odot\Big(\bm{x}_{t}-\Delta t\,\epsilon_{\theta}(\bm{x}_{t},t)+\sigma(t{-}\Delta t)\bm{z}\Big),\end{split}$ (10)
where $\epsilon_{\theta}(\bm{x}_{t},t)$ denotes the predicted noise, $\sigma(t)$ is the noise schedule at step $t$, $\bm{z}\sim\mathcal{N}(0,I)$, and $\Delta t$ is the reverse step size. This keeps unmasked regions faithful while allowing masked regions to be selectively refined.
Holistic Consistency. Even after localized refinement, residual inconsistencies may remain around mask boundaries. To restore smooth transitions, we remove the mask and apply a few global denoising steps, allowing the model to harmonize style and structure across the entire image.
5 Experiments and Analysis
Figure 5: Qualitative results on challenging text-to-image prompts. Compared to Resampling, Best-of-$N$, and Particle Sampling, LoTTS better follows complex prompts. The green borders indicate high-quality generations, red mark lower-quality ones, and brown shows the initial image used as the starting point of localized refinement.
Table 1: Quantitative results on three benchmarks (Pick-a-Pic, DrawBench, COCO2014). Across human-aligned metrics (HPS, AES, Pick, IR) and automated metrics (FID, CLIP), LoTTS achieves the best overall performance across Resampling, Best-of-$N$, and Particle Sampling baselines.